# Improving Generalization and Convergence by Enhancing Implicit Regularization

### v2

@@ Abstract
In this work† Correspondence to: mingzewang@stu.pku.edu.cnleiwu@math.pku.edu.cn, we propose an Implicit Regularization Enhancement (IRE) framework to accelerate the discovery of flat solutions in deep learning, thereby improving generalization and convergence. Specifically, IRE decouples the dynamics of flat and sharp directions, which boosts the sharpness reduction along flat directions while maintaining the training stability in sharp directions. We show that IRE can be practically incorporated with generic base optimizers without introducing significant computational overload. Experiments show that IRE consistently improves the generalization performance for image classification tasks across a variety of benchmark datasets (CIFAR-10/100, ImageNet) and models (ResNets and ViTs). Surprisingly, IRE also achieves a 2×2$\times$2 × speed-up compared to AdamW in the pre-training of Llama models (of sizes ranging from 60M to 229M) on datasets including Wikitext-103, Minipile, and Openwebtext. Moreover, we provide theoretical guarantees, showing that IRE can substantially accelerate the convergence towards flat minima in sharpness-aware minimization (SAM).
@@ 1.1 Related works
Implicit regularization. There have been extensive attempts to explain the mystery of implicit regularization in deep learning (see the survey by Vardi (2023) and references therein). Here, we focus on works related to implicit sharpness regularization. Wu et al. (2018; 2022) and Ma and Ying (2021) provided an explanation of implicit sharpness regularization from a dynamical stability perspective. Moreover, in-depth analysis of SGD dynamics near global minima shows that the SGD noise (Blanc et al., 2020; Li et al., 2022; Ma et al., 2022; Damian et al., 2021) and the edge of stability (EoS)-driven (Wu et al., 2018; Cohen et al., 2021) oscillations (Even et al., 2024) can drive SGD/GD towards flatter minima. Additional studies explored how training components, including learning rate and batch size (Jastrzębski et al., 2017), normalization (Lyu et al., 2022), cyclic LR (Wang and Wu, 2023), influence this sharpness regularization. Furthermore, some works have provided theoretical evidence explaining the superior generalization of flat minima for neural networks (Ma and Ying, 2021; Mulayoff et al., 2021; Wu and Su, 2023; Gatmiry et al., 2023; Wen et al., 2023b). Our work is inspired by this line of research, aiming to boost implicit sharpness regularization by decoupling the dynamics along flat and sharp directions.
@@
Optimizers for large language model (LLM) pre-training. (Momentum) SGD (Sutskever et al., 2013; Nesterov, 1983) and its adaptive variants like Adagrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), and Adam (Kingma and Ba, 2014) have been widely used in DNN training. Despite the efforts in designing better adaptive gradient methods (Liu et al., 2019a; Luo et al., 2019; Heo et al., 2020; Zhuang et al., 2020; Xie et al., 2022b; a), AdamW(Adam+decoupled weight decay) (Loshchilov and Hutter, 2017) has become the default optimizer in LLM pre-training. Recently, Chen et al. (2024) discovered Lion by searching the space of adaptive first-order optimizers; Liu et al. (2024) introduced Sophia, a scalable second-order optimizer. In this paper, we instead empirically demonstrate that IRE can accelerate the convergence of AdamW in the pre-training of Llama models.
@@ 2 An Illustrative Example Motivating IRE
It is worth noting that similar implicit sharpness regularization also holds for SGD (Ma et al., 2022; Li et al., 2022) and SAM (Wen et al., 2023a). In this section, we focus on the above toy model and GD mainly for illustration. In Appendix A, we provide an analogous illustrative analysis of how IRE accelerates the sharpness reduction of SGD. In Section 5, we further provide theoretical evidence to show that IRE can boost the implicit sharpness regularization of SAM.
@@ 4.1.1 Validating our motivation
To show that IRE can accelerate the sharpness reduction, we train WideResNet-16-8 (Zagoruyko and Komodakis, 2016) on CIFAR-10 dataset (Krizhevsky and Hinton, 2009) by SAM-IRE (with $K=10$, varying $\kappa$ and $\gamma$). Here, we incoporate IRE into SAM starting from the $30$-th epochs. We vary $\gamma\in\{0.8,0.9,0.95\}$ and $\kappa\in\{0,2,5,10\}$. Regarding the learning rate (LR), both constant LR and decayed LR are considered. The sharpness is measured by $\operatorname{Tr}(\nabla^{2}\mathcal{L}(\bm{\theta}))$. Further experimental details can be found in Appendix B.
@@ 4.1.2 IRE can consistently improve generalization
Convolutional Neural Networks (CNNs). In this experiment, we first consider the classification of CIFAR-{10,100} with WideResNet-28-10 (Zagoruyko and Komodakis, 2016) and ResNet-56 (He et al., 2016). Both SGD and SAM optimizers are adopted. All the experiments use base data augmentation and label smoothing. For SGD-IRE/SAM-IRE, we fix $K=10$, and tune hyperparameters $\gamma$ and $\kappa$ via a grid search over $\gamma\in\{0.99,0.9,0.8\}$ and $\kappa\in\{1,2\}$. The total epochs are set to 100 for CIFAR-10 and 200 for CIFAR-100, and we switch from SGD/SAM to SGD-IRE/SAM-IRE when the training loss approaches 0.1. The other experimental details are deferred to Appendix B and the results are shown in Table 1.
Secondly, we evaluate IRE for training ResNet-50 on ImageNet (Deng et al., 2009). The experimental details are deferred to Appendix B and the results are shown in Table 3.
Vision Transformers (ViTs). We also examine the impact of IRE on generalization of ViT-T and ViT-S (Dosovitskiy et al., 2020) on CIFAR100. The default optimizers used are AdamW and SAM (Mueller et al., 2024). Strong data augmentations (basic + AutoAugment) are utilized. The total epochs are set to $200$ and we switch from AdamW/SAM to AdmIRE/SAM-IRE when the training loss approaches $0.5$. For AdmIRE/SAM-IRE, we fix $K=10$, and tune hyperparameters $\gamma$ and $\kappa$ via a grid search over $\gamma\in\{0.99,0.9,0.8\}$ and $\kappa\in\{20,50\}$. Other experimental details are deferred to Appendix B. The results are shown in Table 3.
Additionally, we evaluate IRE for training ViT-S on ImageNet. The experimental details are deferred to Appendix B and the results are shown in Table 4.
@@ 4.2.1 Computational efficiency and hyperparameter robustness
More experimental details and results are deferred to Appendix B.
@@ 4.2.2 Experiments on Llama models
Additionally, gradient clipping is adopted to maintain the training stability (Pascanu et al., 2012). First, we tune lr_max in AdamW for each of the three experiments, separately. The optimal lr_max identified for these three experiments is all 6e-4. Then, both AdamW and AdmIRE are trained using this optimal lr_max. For more details, please refer to Appendix B.
AdmIRE is 2×2$\times$2 × faster than AdamW. The results are reported in Figure 4. We can see that AdmIRE consistently achieves a 2.1×2.1$\times$2.1 × speedup compared with well-tuned AdamW for all three cases. However, the underlying mechanism remains unclear and is left for future work.
@@ 5.2 Theoretical results
$\mathbb{E}_{\bm{\xi}_{t}}[\bm{z}_{t+1}]=\bm{z}_{t}-\frac{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}(1+\kappa)}\eta\rho^{2}}{p}\nabla_{\mathcal{M}}\operatorname{Tr}\left[\nabla^{2}\mathcal{L}(\bm{z}_{t})/2\right]+\mathcal{O}(\eta^{3/2}\rho^{2}).$
Note that $\rho=\mathcal{O}(\sqrt{\eta})$ and $\kappa$ can be as large as $1/\rho$. Consequently, the effective LR of minimizing the trace of Hessian can be selected as large as $\eta_{\rm eff}=(\kappa+1)\eta\rho^{2}/p=\mathcal{O}(\eta^{1.5}/p)$. In contrast, that of average SAM is at most $\mathcal{O}(\eta^{2}/p)$. The proof of Theorem 5.5 can be found in Appendix C.
@@ 5.2.2 IRE on standard SAM: An $\Omega(1/\rho)$ acceleration
This subsection delves into IRE on standard SAM (7), which is more widely used and often yields better performance than average SAM (8). However, since standard SAM (7) requires stochastic gradients $\nabla\mathcal{L}_{i}(\bm{\theta})$ ($i\in[n]$), we need an additional assumption regarding the features on the manifold (see Setting D.1), which is commonly used in the literature (Du et al., 2018; 2019; Li et al., 2022; Arora et al., 2022; Wen et al., 2023a). We defer it to Appendix D due to space constraints. Under this Setting, Assumption 5.1 holds naturally with $m=n$.
During Phase I of GF, Definition 5.3 ensures that there exists $t<\infty$ such that $\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)$ for any $\alpha\in[0,1)$. We define $T_{\rm I}$ as the hitting time: $T_{\rm I}:=\inf\{t\geqslant 0:\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)\}$. Then the following result holds for Phase II, whose proof can be founded in Appendix D.
@@
Under Setting D.1, if $\eta,\rho=\mathcal{O}(1)$ in SAM (7), $\kappa\leqslant 1/\rho$, and $\mathcal{P}_{t}=P_{n+1:p}(\nabla^{2}\mathcal{L}(\bm{\theta}_{t}))$ in IRE (4), then with high probability at least $1-T_{\rm II}^{2}\exp\left(-\Omega(1/\eta^{\alpha})\right)$, $\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)$ holds for all $T_{\rm I}\leqslant t\leqslant T$. Moreover, the effective dynamics $\bm{z}_{t}=\Phi(\bm{\theta}_{t})\in\mathcal{M}$ satisfies:
$\mathbb{E}_{i_{t}}[\bm{z}_{t+1}]=\bm{z}_{t}-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}(1+\kappa)}\eta\rho^{2}\nabla_{\mathcal{M}}\operatorname{Tr}\left[\nabla^{2}\mathcal{L}(\bm{z}_{t})/2\right]+\mathcal{O}\left((\kappa+1)\eta\rho^{2}(\rho+\eta^{1-\alpha})\right).$
@@ 6 Conclusion
In this work, we propose a novel IRE framework to enhance the implicit sharpness regularization of base optimizers. Experiments demonstrate that IRE not only consistently improves generalization but also accelerates loss convergence in the pre-training of Llama models of various sizes. For future work, there are two urgent directions: 1) understanding why IRE can accelerate convergence, which may require studying the interplay between IRE and the Edge of Stability (EoS) (Wu et al., 2018; Jastrzębski et al., 2017; Cohen et al., 2021); and 2) conducting a larger-scale investigation into the acceleration of AdmIRE compared to AdamW in LLM pre-training, as well as the downstream performance of the LLMs pre-trained by AdmIRE.
@@ Acknowledgments
Lei Wu is supported by the National Key R&D Program of China (No. 2022YFA1008200) and National Natural Science Foundation of China (No. 12288101). Mingze Wang is supported in part by the National Key Basic Research Program of China (No. 2015CB856000). We thank Dr. Hongkang Yang, Liu Ziyin, Liming Liu, Zehao Lin, Hao Wu, and Kai Chen for helpful discussions.
@@
Appendix A Proofs for SGD in Section 2

### v3

@@ Abstract
In this work, we propose an Implicit Regularization Enhancement (IRE) framework to accelerate the discovery of flat solutions in deep learning, thereby improving generalization and convergence. Specifically, IRE decouples the dynamics of flat and sharp directions, which boosts the sharpness reduction along flat directions while maintaining the training stability in sharp directions. We show that IRE can be practically incorporated with generic base optimizers without introducing significant computational overload. Experiments show that IRE consistently improves the generalization performance for image classification tasks across a variety of benchmark datasets (CIFAR-10/100, ImageNet) and models (ResNets and ViTs). Surprisingly, IRE also achieves a 2×2$\times$2 × speed-up compared to AdamW in the pre-training of Llama models (of sizes ranging from 60M to 229M) on datasets including Wikitext-103, Minipile, and Openwebtext. Moreover, we provide theoretical guarantees, showing that IRE can substantially accelerate the convergence towards flat minima in sharpness-aware minimization (SAM).
† Correspondence to: Mingze Wang and Lei Wu.
@@ 1.1 Related works
Implicit sharpness regularization. There have been extensive attempts to explain the mystery of implicit regularization in deep learning (see the survey by Vardi (2023) and references therein). Here, we focus on works related to implicit sharpness regularization. Wu et al. (2018, 2022) and Ma and Ying (2021) provided an explanation of implicit sharpness regularization from a dynamical stability perspective. Moreover, in-depth analysis of SGD dynamics near global minima shows that the SGD noise (Blanc et al., 2020; Li et al., 2022; Ma et al., 2022; Damian et al., 2021) and the edge of stability (EoS)-driven (Wu et al., 2018; Cohen et al., 2021) oscillations (Even et al., 2024) can drive SGD/GD towards flatter minima. Additional studies explored how training components, including learning rate and batch size (Jastrzębski et al., 2017), normalization (Lyu et al., 2022), cyclic LR (Wang and Wu, 2023), influence this sharpness regularization. Furthermore, some works have provided theoretical evidence explaining the superior generalization of flat minima for neural networks (Ma and Ying, 2021; Mulayoff et al., 2021; Wu and Su, 2023; Gatmiry et al., 2023; Wen et al., 2023b). Our work is inspired by this line of research, aiming to boost implicit sharpness regularization by decoupling the dynamics along flat and sharp directions.
@@
Optimizers for large language model (LLM) pre-training. (Momentum) SGD (Sutskever et al., 2013; Nesterov, 1983) and its adaptive variants like Adagrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), and Adam (Kingma and Ba, 2014) have been widely used in DNN training. Despite the efforts in designing better adaptive gradient methods (Liu et al., 2019a; Luo et al., 2019; Heo et al., 2020; Zhuang et al., 2020; Xie et al., 2022b, a), AdamW(Adam+decoupled weight decay) (Loshchilov and Hutter, 2017) has become the default optimizer in LLM pre-training. Recently, Chen et al. (2024) discovered Lion by searching the space of adaptive first-order optimizers; Liu et al. (2024) introduced Sophia, a scalable second-order optimizer. In this paper, we instead empirically demonstrate that IRE can accelerate the convergence of AdamW in the pre-training of Llama models.
@@ 2 An Illustrative Example Motivating IRE
It is worth noting that similar implicit sharpness regularization also holds for SGD (Ma et al., 2022; Li et al., 2022) and SAM (Wen et al., 2023a). In this section, we focus on the above toy model and GD mainly for illustration. In Appendix B, we provide an analogous illustrative analysis of how IRE accelerates the sharpness reduction of SGD. In Section 5, we further provide theoretical evidence to show that IRE can boost the implicit sharpness regularization of SAM.
@@ 4.1.1 Validating our motivation
To show that IRE can accelerate the sharpness reduction, we train WideResNet-16-8 (Zagoruyko and Komodakis, 2016) on CIFAR-10 dataset (Krizhevsky and Hinton, 2009) by SAM-IRE (with $K=10$, varying $\kappa$ and $\gamma$). Here, we incoporate IRE into SAM starting from the $30$-th epochs. We vary $\gamma\in\{0.8,0.9,0.95\}$ and $\kappa\in\{0,2,5,10\}$. Regarding the learning rate (LR), both constant LR and decayed LR are considered. The sharpness is measured by $\operatorname{Tr}(\nabla^{2}\mathcal{L}(\bm{\theta}))$. Further experimental details can be found in Appendix C.
@@ 4.1.2 IRE can consistently improve generalization
Convolutional Neural Networks (CNNs). In this experiment, we first consider the classification of CIFAR-{10,100} with WideResNet-28-10 (Zagoruyko and Komodakis, 2016) and ResNet-56 (He et al., 2016). Both SGD and SAM optimizers are adopted. All the experiments use base data augmentation and label smoothing. For SGD-IRE/SAM-IRE, we fix $K=10$, and tune hyperparameters $\gamma$ and $\kappa$ via a grid search over $\gamma\in\{0.99,0.9,0.8\}$ and $\kappa\in\{1,2\}$. The total epochs are set to 100 for CIFAR-10 and 200 for CIFAR-100, and we switch from SGD/SAM to SGD-IRE/SAM-IRE when the training loss approaches 0.1. The other experimental details are deferred to Appendix C and the results are shown in Table 1.
Secondly, we evaluate IRE for training ResNet-50 on ImageNet (Deng et al., 2009). The experimental details are deferred to Appendix C and the results are shown in Table 3.
Vision Transformers (ViTs). We also examine the impact of IRE on generalization of ViT-T and ViT-S (Dosovitskiy et al., 2020) on CIFAR100. The default optimizers used are AdamW and SAM (Mueller et al., 2024). Strong data augmentations (basic + AutoAugment) are utilized. The total epochs are set to $200$ and we switch from AdamW/SAM to AdmIRE/SAM-IRE when the training loss approaches $0.5$. For AdmIRE/SAM-IRE, we fix $K=10$, and tune hyperparameters $\gamma$ and $\kappa$ via a grid search over $\gamma\in\{0.99,0.9,0.8\}$ and $\kappa\in\{20,50\}$. Other experimental details are deferred to Appendix C. The results are shown in Table 3.
Additionally, we evaluate IRE for training ViT-S on ImageNet. The experimental details are deferred to Appendix C and the results are shown in Table 4.
@@ 4.2.1 Computational efficiency and hyperparameter robustness
More experimental details and results are deferred to Appendix C.
@@ 4.2.2 Experiments on Llama models
Additionally, gradient clipping is adopted to maintain the training stability (Pascanu et al., 2012). First, we tune lr_max in AdamW for each of the three experiments, separately. The optimal lr_max identified for these three experiments is all 6e-4. Then, both AdamW and AdmIRE are trained using this optimal lr_max. For more details, please refer to Appendix C.
AdmIRE is 2×2$\times$2 × faster than AdamW. The results are reported in Figure 4. We can see that AdmIRE consistently achieves a 2.1×2.1$\times$2.1 × speedup compared with well-tuned AdamW for all three cases.
@@
Notice that the primary motivation behind IRE is to speed up the sharpness reduction, which only requires to increase learning rate along completely flat (zero-curvature) directions. However, practical implementation may also increase the learning rate along directions with small but non-zero curvatures, which can further speed up loss convergence. A thorough explanation for the significant acceleration provided by this approach is left for future research.
Table 6: Comparison of the sharpness of the solutions found by AdamW/AdmIRE.
AdamW AdmIRE
training steps 100k 50k
final $\mathcal{L}(\bm{\theta})$ 2.47 2.47
final $\operatorname{Tr}(\nabla^{2}\mathcal{L}(\bm{\theta}))$ 120.41 88.86
We further assess the sharpness reduction capability of IRE for LLM pre-training. Specifically, we compare the sharpness of solutions, $\operatorname{Tr}(\nabla^{2}\mathcal{L}(\bm{\theta}))$, found by AdamW/AdmIRE during pre-training of Llama (60M) on wiki-103 dataset (corresponding to Figure 4 (left)). The results shown in Table 6 demonstrate that AdamIRE not only achieves the same loss in only half the iterations required by AdamW, but also the solutions found by AdmIRE are significantly flatter than that found by AdamW.
Recently, Liu et al. (2023) revealed a strong correlation between the sharpness and downstream task performance, suggesting that for models with the same pre-training loss, flatter solutions yield better performance on downstream tasks. Based on this, we hypothesize that the solutions found by IRE may also have better performance in downstream tasks, which we leave to future work.
@@ 5.2 Theoretical results
$\mathbb{E}_{\bm{\xi}_{t}}[\bm{z}_{t+1}]=\bm{z}_{t}-\frac{{\color[rgb]{.5,0,.5}(1+\kappa)}\eta\rho^{2}}{p}\nabla_{\mathcal{M}}\operatorname{Tr}\left[\nabla^{2}\mathcal{L}(\bm{z}_{t})/2\right]+\mathcal{O}(\eta^{3/2}\rho^{2}).$
Note that $\rho=\mathcal{O}(\sqrt{\eta})$ and $\kappa$ can be as large as $1/\rho$. Consequently, the effective LR of minimizing the trace of Hessian can be selected as large as $\eta_{\rm eff}=(\kappa+1)\eta\rho^{2}/p=\mathcal{O}(\eta^{1.5}/p)$. In contrast, that of average SAM is at most $\mathcal{O}(\eta^{2}/p)$. The proof of Theorem 5.5 can be found in Appendix D.
@@ 5.2.2 IRE on standard SAM: An $\Omega(1/\rho)$ acceleration
This subsection delves into IRE on standard SAM (7), which is more widely used and often yields better performance than average SAM (8). However, since standard SAM (7) requires stochastic gradients $\nabla\mathcal{L}_{i}(\bm{\theta})$ ($i\in[n]$), we need an additional assumption regarding the features on the manifold (see Setting E.1), which is commonly used in the literature (Du et al., 2018, 2019; Li et al., 2022; Arora et al., 2022; Wen et al., 2023a). We defer it to Appendix E due to space constraints. Under this Setting, Assumption 5.1 holds naturally with $m=n$.
During Phase I of GF, Definition 5.3 ensures that there exists $t<\infty$ such that $\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)$ for any $\alpha\in[0,1)$. We define $T_{\rm I}$ as the hitting time: $T_{\rm I}:=\inf\{t\geqslant 0:\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)\}$. Then the following result holds for Phase II, whose proof can be founded in Appendix E.
@@
Under Setting E.1, if $\eta,\rho=\mathcal{O}(1)$ in SAM (7), $\kappa\leqslant 1/\rho$, and $\mathcal{P}_{t}=P_{n+1:p}(\nabla^{2}\mathcal{L}(\bm{\theta}_{t}))$ in IRE (4), then with high probability at least $1-T_{\rm II}^{2}\exp\left(-\Omega(1/\eta^{\alpha})\right)$, $\left\|\bm{\theta}_{t}-\Phi(\bm{\theta}_{t})\right\|=\mathcal{O}(\eta^{1-\alpha}\rho)$ holds for all $T_{\rm I}\leqslant t\leqslant T$. Moreover, the effective dynamics $\bm{z}_{t}=\Phi(\bm{\theta}_{t})\in\mathcal{M}$ satisfies:
$\mathbb{E}_{i_{t}}[\bm{z}_{t+1}]=\bm{z}_{t}-{\color[rgb]{.5,0,.5}(1+\kappa)}\eta\rho^{2}\nabla_{\mathcal{M}}\operatorname{Tr}\left[\nabla^{2}\mathcal{L}(\bm{z}_{t})/2\right]+\mathcal{O}\left((\kappa+1)\eta\rho^{2}(\rho+\eta^{1-\alpha})\right).$
@@ 6 Conclusion
In this work, we propose a novel IRE framework to enhance the implicit sharpness regularization of base optimizers. Experiments demonstrate that IRE not only consistently improves generalization but also accelerates loss convergence in the pre-training of Llama models of various sizes. The code is available at https://github.com/wmz9/IRE.
For future work, there are two urgent directions: 1) understanding why IRE can accelerate convergence, which may require studying the interplay between IRE and the Edge of Stability (EoS) (Wu et al., 2018; Jastrzębski et al., 2017; Cohen et al., 2021); and 2) conducting a larger-scale investigation into the acceleration of AdmIRE compared to AdamW in LLM pre-training, as well as the downstream performance of the LLMs pre-trained by AdmIRE.
@@ Acknowledgments
Lei Wu is supported by the National Key R&D Program of China (No. 2022YFA1008200) and National Natural Science Foundation of China (No. 12288101). Mingze Wang is supported in part by the National Key Basic Research Program of China (No. 2015CB856000). We thank Dr. Hongkang Yang, Liu Ziyin, Liming Liu, Zehao Lin, Hao Wu, and Kai Chen for helpful discussions and anonymous reviewers for their valuable suggestions.
@@
Appendix A Other Related Works
Other Implicit Biases. Beyond the implicit sharpness regularization, numerous other attempts have explored implicit biases in deep learning algorithms Vardi (2023). Among these, a popular research line is the max-margin bias, which suggests that optimizers favor the solutions with large margin, which generalizes well. Soudry et al. (2018) showed that GD converges to max-margin solutions under exponentially-tailed loss on linearly separable data, albeit with an extremely slow rate $\mathcal{O}(1/\log t)$. Furthermore, Nacson et al. (2019c) studied this bias for SGD, Ji and Telgarsky (2019b) investigated linearly non-separable data, and Ji et al. (2020) analyzed the effects of the tail behavior of loss functions.
To achieve faster margin maximization, Nacson et al. (2019b); Ji and Telgarsky (2021) demonstrated that GD with aggressively loss-scaled step sizes can achieve a faster polynomial rate of $\mathcal{O}(1/t)$. Building on this, Ji et al. (2021); Wang et al. (2022) proposed momentum-based gradient methods which achieve a rate of $\mathcal{O}(1/t^{2})$. Recently, Wang et al. (2024) established that the polynomial rates for most previous algorithms are tight, and proposed a progressive rescaling gradient descent method that achieves margin maximization exponentially fast $\mathcal{O}(e^{-\Omega(t)})$.
The margin-maximization bias has also been studied for nonlinear models.  Ji and Telgarsky (2019a); Gunasekar et al. (2018) examined deep linear networks, while Chizat and Bach (2020) focused on wide two-layer ReLU networks. Notably, Nacson et al. (2019a); Lyu and Li (2019); Ji and Telgarsky (2020) demonstrated that for general homogeneous networks, Gradient Flow (GF) and GD converge to solutions corresponding the KKT point of the max-margin problem. Recently, Kunin et al. (2023) extended this analysis to quasi-homogeneous networks. For two-layer (leaky-)ReLU neural networks, Lyu et al. (2021); Vardi et al. (2022); Wang and Ma (2023) examined whether the convergent KKT point of GF correspond to global optima of the max-margin problem.
Future work could investigate whether the IRE framework can also enhance the margin maximization bias, although it is primarily designed for enhancing the implicit sharpness regularization.
Additionally, Woodworth et al. (2020); Pesme et al. (2021); Nacson et al. (2022); Pesme and Flammarion (2023); Even et al. (2023) conducted fine-grained analyses of training dynamics, examining how initialization and step size impact (S)GD’s minima selection in linear diagonal networks.
Appendix B Proofs for SGD in Section 2